OpenStudy (anonymous):

Zoey opened a savings account four years ago with a deposit of $935.15. The interest on the account compounds monthly. If Zoey’s current account balance is $1,128.16, what is the interest rate on the account? 0.4% 4.7% 14.2% 18.9%

OpenStudy (anonymous):

@jdoe0001 can u help on this one?

OpenStudy (anonymous):

Don't want the answer I wanna know how to find it? :)

OpenStudy (jdoe0001):

sorry the system is a bit lagged.... you know the compound interest equation, right?

OpenStudy (jdoe0001):

anyhow, the compound interest balance, can be found by \(\bf b = balance\\ p = principal\\ r = rate\\ n = periods\\ t = years\\ \quad \\ \quad \\ b = p\left(1+\cfrac{r}{n}\right)^{nt}\)

OpenStudy (jdoe0001):

so you need the rate, that is, you need "r" so let's solve for "r"

OpenStudy (jdoe0001):

\(\bf b = balance\\ p = principal\\ r = rate\\ n = periods\\ t = years\\ \quad \\ \quad \\ b = p\left(1+\cfrac{r}{n}\right)^{nt}\\\\ \quad \\ \cfrac{b}{p} = \left(1+\cfrac{r}{n}\right)^{nt}\\ \textit{taking "nt" root from both sides}\\ \large {\sqrt[nt]{\cfrac{b}{p}} = 1+\cfrac{r}{n} \implies \sqrt[nt]{\cfrac{b}{p}}-1 = \cfrac{r}{n}\\ \quad \implies n\cdot \left(\sqrt[nt]{\cfrac{b}{p}}-1\right) = r}\)